# The Expressive Power of Transformers with Chain of Thought

### 20 Mar 2024 v4

@@ 4 Upper Bounds for Transformer Decoders
Having shown lower bounds on transformers with $t(n)$ steps, we present two different upper bounds: one that relates transformer decoders to time complexity classes, and one that relates them to space complexity classes. The relative strength of the two different bounds will vary depending on $t(n)$.
A simple upper bound on transformers with chain of thought can be obtained based on the fact that transformers can be simulated using a quadratic number of arithmetic operations.
@@
$\subseteq\mathsf{SPACE}(t(n)+c\log n+c\log t(n))$
$=\mathsf{SPACE}(t(n)+\log n).\qed$
@@ Acknowledgements
We thank David Chiang for identifying a mismatch between the transformer definition in an earlier version of this paper and standard pre-norm transformers. We also appreciate helpful comments from Gabriel Faria, Ofir Press, Abulhair Saparov, Jason Wei, and Avi Wigderson, as well as researchers in ML2 at NYU and at AI2. WM was supported by NSF Award 1922658, an NSF Graduate Research Fellowship, and AI2.
@@ Appendix A Transformer Components
This section recalls the definition from Merrill & Sabharwal (2023a) for the components of the transformer layer. We assume a pre-norm (Xiong et al., 2020) parameterization of the transformer for concreteness and because this is more standard in newer transformers. As stated in Section 2.1, we also allow multiple pre-norms of different projections of the hidden state for our lower-bound constructions. Concretely, we define $\mathsf{multi\_layer\_norm}(\mathbf{v})$ as follows.
@@
Let $m$ be divisible by $k$. For $1\leq i\leq k$, let $\mathbf{W}_{i}:m\to m/k$ be a trainable parameter matrix that projects $m$ dimensional vectors to $m/k$ dimensional vectors. Let $\langle\cdot\rangle_{i=1}^{k}$ denote iterated concatenation. The multi-pre-norm of $\mathbf{v}\in\mathbb{R}^{m}$ is defined as
$\mathsf{multi\_layer\_norm}(\mathbf{v})=\left\langle\mathsf{layer\_norm}(\mathbf{W}_{i}\mathbf{v})\right\rangle_{i=1}^{k}.$
A.1 Transformer Embeddings
@@ A.3 Activation Block
Appendix B Layer-Norm Hash
@@
Another useful property of the layer-norm hash is the following:

### 11 Apr 2024 v5

@@ 4 Upper Bounds for Transformer Decoders
Having shown lower bounds on transformers with $t(n)$ steps, we present two different upper bounds: one that relates transformer decoders to time complexity classes, and one that relates them to space complexity classes. The relative strength of the two different bounds will vary depending on $t(n)$. A simple upper bound on transformers with chain of thought can be obtained based on the fact that transformers can be simulated using a quadratic number of arithmetic operations.
@@
$\subseteq\mathsf{SPACE}(t(n)+\log n).\qed$
@@ Acknowledgements
We thank David Chiang for the valuable feedback and for identifying a mismatch between the transformer definition in an earlier version of this paper and standard pre-norm transformers. We also appreciate helpful comments from Gabriel Faria, Ofir Press, Abulhair Saparov, Jason Wei, and Avi Wigderson, as well as researchers in ML2 at NYU and at AI2. WM was supported by NSF Award 1922658, an NSF Graduate Research Fellowship, and AI2.
@@ Appendix A Transformer Components
This section formally defines our generalizations of pre-norm and then recalls the definition from Merrill & Sabharwal (2023a) for the components of the transformer layer.
A.1 Generalized Pre-Norm
We assume a pre-norm (Xiong et al., 2020) parameterization of the transformer for concreteness, as this is more standard in newer transformers. As stated in Section 2.1, we allow transformer sublayers to apply a linear projection before layer-norm. Concretely, we define $\mathsf{proj\_layer\_norm}(\mathbf{v})$ as follows:
Definition 4 (Projected pre-norm).
Let $\mathbf{M}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ be a parameter matrix that projects $m$-dimensional vectors to $m$-dimensional vectors.
$\mathsf{proj\_layer\_norm}(\mathbf{v};\mathbf{M})=\mathsf{layer\_norm}(\mathbf{M}\mathbf{v}).$
We will omit the parameter $\mathbf{M}$ for convenience, instead writing $\mathsf{proj\_layer\_norm}(\mathbf{v})$.
Our proofs also allow multiple pre-norms of different projections of the hidden state for our lower-bound constructions. Concretely, we define $\mathsf{multi\_layer\_norm}(\mathbf{v})$ as follows.
@@ A.1 Generalized Pre-Norm
Let $m$ be divisible by $k$. For $1\leq i\leq k$, let $\mathbf{M}_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m/k}$ be a parameter matrix that projects $m$-dimensional vectors to $m/k$-dimensional vectors. Let $\langle\cdot\rangle_{i=1}^{k}$ denote iterated concatenation. The multi-pre-norm of $\mathbf{v}\in\mathbb{R}^{m}$ is defined as
$\mathsf{multi\_layer\_norm}(\mathbf{v};\mathbf{M}_{1},\ldots\mathbf{M}_{k})=\left\langle\mathsf{proj\_layer\_norm}(\mathbf{v};\mathbf{M}_{i})\right\rangle_{i=1}^{k}.$
As for projected pre-norm, we will omit the parameters $\mathbf{M}_{1},\ldots\mathbf{M}_{k}$ for multi-pre-norm, instead writing $\mathsf{multi\_layer\_norm}(\mathbf{v})$.
As noted earlier, projected pre-norm can simulate multi-pre-norm:
See 1
Proof.
We will simulate a multi-pre-norm layer that takes as input
$\mathsf{layer\_norm}(\mathbf{M}_{1}\mathbf{v}),\ldots,\mathsf{layer\_norm}(\mathbf{M}_{k}\mathbf{v})$
using only projected pre-norm layers. The idea is to use $k$ different projected pre-norm layers (one for each input layer norm). At layer $i$, the layer takes as input $\mathsf{layer\_norm}(\mathbf{M}_{i}\mathbf{v})$ and writes this to the residual stream. Then, after these $k$ layers, a final layer reads as input
$\mathsf{layer\_norm}\left(\left\langle\mathsf{layer\_norm}(\mathbf{M}_{i}\mathbf{v})\right\rangle_{i=1}^{k}\right).$
Since each vector in the concatenation has unit norm, this is equivalent to
$\frac{1}{\sqrt{k}}\left\langle\mathsf{layer\_norm}(\mathbf{M}_{i}\mathbf{v})\right\rangle_{i=1}^{k}.$
It follows that this layer receives essentially the same input as the original multi-pre-norm layer, up to a constant factor. The original weights for the layer can be multiplied by $\sqrt{k}$ to implement the same computation as the original layer. To make sure that $\sqrt{k}$ is exactly representable, we can pad the number of entries so that $k$ is a perfect square. ∎
A.2 Transformer Embeddings
@@ A.4 Activation Block
Appendix B Turing MAchines
A Turing machine takes as input a string $\sigma\in\Sigma^{*}$. A configuration of a Turing machine is a finite state $q$ along with the contents of an input tape $c^{0}$, $k$ work tapes $c^{1},\ldots,c^{k}$, and an output tape $c^{k+1}$. Finally, for each tape $\tau$, a configuration specifies a head position $h^{\tau}$. We start with the initial state $q_{0}$ and the input tape $c^{0}_{0}$ containing $\sigma$ starting at position 00 with infinite $b$’s on each side, and $h^{0}_{0}=0$. All other tapes start containing all $b$’s and with their head at 00. At each time step $i$, if $q_{i}\not\in F$, we recurrently update the configuration by first computing:
Appendix C Layer-Norm Hash
@@
Appendix D Input Tape Retrieval via the Layer-Norm Hash
We now describe an attention head that uses the layer-norm hash to read from the input tape in Theorem 2. Define a sequence $h_{1},\ldots,h_{i-1}$, which represents Turing machine tape position in Theorem 2.
We define the following layer-norm hash based quantity, which is instantiated in Theorem 2 in a particular way:
$\phi_{i}=\begin{cases}\phi(i,1)&\textrm{if}\;1\leq i\leq n\\ \phi(h_{i},1)&\textrm{otherwise}.\end{cases}$
The attention head we construct can then be described as follows:
  • Query: $\langle\phi_{i},-1\rangle$
  • Key: $\langle\phi_{j},\mathbbm{1}[n<j]\rangle$
  • Value: $\mathbf{v}_{j}\triangleq\langle\phi_{j},\sigma_{j}\rangle$
Let $\mathbf{\bar{v}}\triangleq\langle\bar{\phi},\bar{\sigma}\rangle$ be the head output. This head obeys the following properties:
Lemma 3.
Let $i>n$. Then, $1\leq h_{i}\leq n$ if and only if $\bar{\phi}=\phi_{i}$.
Proof.
We proceed by considering the two directions.
Forward Direction. The query-key inner product has two terms $\kappa^{1}_{ij}+\kappa^{2}_{ij}$. By Lemma 2, $\kappa^{1}_{ij}$ is maximized either when $h^{0}_{i}=j$ (and $1\leq j\leq n$) or when $h_{i}=h_{j}$ (and $n<j$). However, if $n<j$, the second term $\kappa^{2}_{ij}=1$. Thus, the attention score is maximized uniquely when $j=h_{i}$, so the head retrieves $\mathbf{\bar{v}}=\langle\phi(h_{i},1),\sigma_{h_{i}}\rangle$. Thus, $\bar{\phi}=\phi(h_{1},1)=\phi_{i}$.
Backward Direction. We establish bidirectionality by proving the contrapositive. Assume that either $h_{i}<1$ or $h_{i}>n$. The head retrieves $\bar{\phi}=\frac{1}{\left\lvert M\right\rvert}\sum_{j\in M}\phi(j,1)$ for some $M\subseteq\{1,\ldots,n\}$. It holds that, for all $1\leq j\leq n$, $h_{i}<j$, or the other way around (i.e., for all $1\leq j\leq n$, $h_{i}>j$). Thus, by Lemma 5, $\bar{\phi}\neq\phi(h_{i},1)=\phi_{i}$. ∎
The following property also emerges from the proof of the forward direction in Lemma 3:
Lemma 4.
Let $i>n$. Then, if $1\leq h_{i}\leq n$, $\bar{\sigma}=\sigma_{h_{i}}$.
The backward direction in Lemma 3 relies on the following lemma:
@@ Appendix D Input Tape Retrieval via the Layer-Norm Hash
Appendix E Rightmost Retrieval via the Layer-Norm Hash
We now describe an attention head that can attend to the rightmost token satisfying a certain property, capturing the construction in Theorem 2 to retrieve the most recent write to a Turing machine work tape. Define a sequence $h_{1},\ldots,h_{i-1}$, which represents Turing machine tape position in Theorem 2. As is natural for Turing machine tapes, we assume that if $h\neq h_{i}$ for all $i$, then it must be that $h\prec h_{i}$ for all $i$, where $\prec$ is fixed as either $>$ or $<$.
Let $f(i)$ be a tie-breaking term that we will define later in Section E.1. We define two layer-norm hash quantities:
$\phi_{i}$ $\triangleq\phi(h_{1}/i,1/i)$
$\psi_{i}$ $\triangleq\phi(f(i),1).$
Recall that $e_{1}=\langle 1,0,0,0\rangle$. Construct an attention head as follows:
  • Query: $\langle\phi_{i},e_{1}\rangle$
  • Key: $\langle\phi_{j},-\psi_{j}\rangle$
  • Value: $\mathbf{v}_{j}\triangleq\langle\phi_{j},\delta_{j}\rangle$
Let $\mathbf{\bar{v}}\triangleq\langle\bar{\phi},\bar{\delta}\rangle$ be the head output. The following properties hold for such a head:
Lemma 6.
There exists $j<i$ such that $h_{i}=h_{j}$ if and only if $\bar{\phi}=\phi_{i}$.
Proof.
We proceed by considering the two directions.
Forward Direction. The query-key inner product has two terms $\kappa^{1}_{ij}+\kappa^{2}_{ij}$. By Lemma 2, the first term $\kappa^{1}_{ij}$ is maximized at 1 for each $j$ such that $h_{i}=h_{j}$. For $h_{i}\neq h_{j}$, $\kappa^{1}_{ij}<1-1/(2i^{4})$ by Lemma 8. The second component $\kappa^{2}_{ij}$ monotonically increases with $j$ and satisfies $\kappa^{2}_{ij}<f(i)<1/(2i^{4})$ by Lemma 10. Thus, the attention score is maximized for the largest $j<i$ such that $h^{\tau}_{i}=h^{\tau}_{j}$. Thus, $\mathbf{\bar{v}}=\mathbf{v}_{j}$ and $\bar{\phi}=\phi_{j}$ for this $j$, which means $\bar{\phi}=\phi_{i}$.
Backward Direction. We establish bidirectionality by proving the contrapositive. Assume there is no $j<i$ such that $h_{i}=h_{j}$. Then $\bar{\phi}=\frac{1}{\left\lvert M\right\rvert}\sum_{j\in M}\phi(h_{j},1)$ for some $M\subseteq\{1,\ldots,n\}$. By assumption (top of Appendix E), we have $h_{j}\prec h_{i}$ for all $j<i$. It follows from Lemma 5 that $\bar{\phi}\neq\phi(h_{i},1)=\phi_{i}$, completing the proof. ∎
The following property also emerges from the forward direction of the proof above:
Lemma 7.
If there exists $j<i$ such that $h_{i}=h_{j}$, then $\bar{\delta}=\delta_{j}$ for the greatest such $j$.
E.1 Tie-Breaking Term
The construction above uses a tie-breaker that favors retrieving tokens further to the right. We will justify the construction of such a tie-breaking term here. To begin, we will establish a bound on the layer-norm hash inner product similarity for inexact matches.
Lemma 8.
For any $i\geq 2$, $\phi(i,1)\cdot\phi(i-1,1)\leq 1-1/(2i^{4})$.
Proof.
Consider the squared dot product:
$\big{(}\phi(i,1)\cdot\phi(i-1,1)\big{)}^{2}$ $=\frac{\big{(}\langle i,1,-i,-1\rangle\cdot\langle i-1,1,-(i-1),-1\rangle\big{)}^{2}}{(2i^{2}+2)(2(i-1)^{2}+2)}$
$=\frac{\big{(}i(i-1)+1\big{)}^{2}}{(i^{2}+1)((i-1)^{2}+1)}$
$=\frac{i^{2}(i-1)^{2}+2i(i-1)+1}{i^{2}(i-1)^{2}+i^{2}+(i-1)^{2}+1}$
$=\frac{i^{2}(i-1)^{2}+2i(i-1)+1}{i^{2}(i-1)^{2}+2i(i-1)+2}$
$=1-\frac{1}{i^{2}(i-1)^{2}+2i(i-1)+2}$
$=1-\frac{1}{i^{4}-2i^{3}+3i^{2}-2i+2}$
Since $(1-y/2)^{2}\geq 1-y$ for any $y$, we have $\sqrt{1-y}\leq 1-y/2$ for any $y\leq 1$. Applying this to the right hand side of the above equation, we obtain:
$\phi(i,1)\cdot\phi(i-1,1)$ $=\sqrt{1-\frac{1}{i^{4}-2i^{3}+3i^{2}-2i+2}}$
$\leq 1-\frac{1}{2i^{4}-4i^{3}+6i^{2}-4i+4}$
$\leq 1-\frac{1}{2i^{4}}\ \ \ \text{for}\;i\geq 2\text{,}$
which completes the proof. ∎
To break ties in attention, we aim to construct a function of $i$ that is computable in the transformer, monotonically decreasing with $i$, and smaller than $1/(2i^{4})$. The following definition will accomplish this:
Definition 6.
We define the following inductively:
$f(i,0)$ $=1/i$
$f(i,k+1)$ $=f(i-1,k)-f(i,k).$
By construction, $f(i,k)$ is monotonically increasing and a linear combination of $1/i,1/(i-1),\ldots,1/(i-k)$. The latter property means it is computable by a single multihead self-attention layer. To do this, we construct $k$ heads in parallel, where head $h$ attends to all tokens besides the first $h$ and puts value $1$ at token $h+1$ and 00 elsewhere. [^10] This head can be implemented by setting a flag in the previous layer at each $i$ for whether $i\leq h$ by hardcoding a comparison between $\phi(1,1/i)$ and $\phi(h,1)$. Head $h$ thus computes $1/(i-h)$. We use the linear transformation at the end of the layer to compute $f(i,k)$ via a linear combination of the head outputs.
Lemma 9.
For any $k$ and $i>k$, we have
$f(i,k)=\frac{k!}{\prod_{j=0}^{k}(i-j)}.$
Proof.
By induction over $k$.
Base Case: $i=0$. We have $f(i,0)=0!/i$.
Inductive Case. We analyze the form of $f(i,k+1)$:
$f(i,k+1)$ $=f(i-1,k)-f(i,k)$
$=\frac{k!}{\prod_{j=0}^{k}(i-1-j)}-\frac{k!}{\prod_{j=0}^{k}(i-j)}$ (Inductive assumption)
$=k!\cdot\frac{\prod_{j=0}^{k}(i-j)-\prod_{j=0}^{k}(i-1-j)}{i\prod_{j=1}^{k}(i-j)^{2}(i-k-1)}$ (Form common denominator)
$=k!\cdot\frac{i\prod_{j=1}^{k}(i-j)-(i-k-1)\prod_{j=1}^{k}(i-j)}{i\prod_{j=1}^{k}(i-j)^{2}(i-k-1)}$ (Pull out factors)
$=k!\cdot\frac{(k+1)\prod_{j=1}^{k}(i-j)}{i\prod_{j=1}^{k}(i-j)^{2}(i-k-1)}$ (Distributive property)
$=\frac{(k+1)!}{i\prod_{j=1}^{k}(i-j)(i-k-1)}$ (Simplify)
$=\frac{(k+1)!}{\prod_{j=0}^{k+1}(i-j)}.\qed$
It remains to be shown that $f(i,k)$ can be made smaller than $1/(2i^{4})$. To handle edge cases around small values of $i$, we define:
Definition 7.
Let ${\epsilon}=10^{-10}$. For $i\geq 1$, let
$f(i)=\begin{cases}1/1000-{\epsilon}i&\text{if\ }\;i\leq 4\\ f(i,3)/100&\text{if\ }i\geq 5\end{cases}$
Lemma 10.
For $i\geq 1$, $f(i)<1/(2i^{4})$
Proof.
When $i\leq 4$, we have $f(i)<1/1000<1/(2i^{4})$.
When $i\geq 5$, by Lemma 9, we have:
$\frac{f(i,3)}{100}$ $=\frac{3!}{100i(i-1)(i-2)(i-3)}$
It can be verified that the values of $i$ for which this expression equals $1/(2i^{4})$ are all in the interval $[0,5)$, and that for $i\geq 5$, $(100/6)i(i-1)(i-2)(i-3)>2i^{4}$. ∎